A private airplane leaves an airport and flies due south at 192 km/h. Two hours later a jet leaves the same airport and flies due south at 960 km/h. When will the jet overtake the plane?
Answer by mananth(16948)

the distance traveled by both is same when one catches up with the other.
Let them meet at distance x
Plane speed= 192 mph , starts at
Jet speed= 960 mph starts at

Time difference between the starting of the two= 2.00 hours

Time taken by Plane = X / 192
Time taken by Jet = X / 960

X / 192 - X / 960 = 2

LCD - 960
Multiply equation by 960 960

5 x - 1 x = 1920
4 x = 1920
/ 4
x= 480 miles
Jet speed= 960 mph
Distance = 480 miles
Time taken = 480 / 960
Time taken = 0.5 hours